OpenStudy (anonymous):

What is the least positive integer k for which 15k is the cube of a number?

OpenStudy (mayankdevnani):

If k = 225 then 15k = 3375 = 15^3 There cannot be a smaller one. Since 15 = 3*5 which are both prime numbers, the cube must be (3^3)*(5^3) = 3375. http://answers.yahoo.com/question/index?qid=20080916035020AAsBCsE

OpenStudy (mayankdevnani):

ok @Areesha.1D

OpenStudy (anonymous):

No i still dont understand. What is the rule or principle fr basic understanding of this?

OpenStudy (anonymous):